Determine the Laplace transform of
f(t)=t^2e^2t
1
Expert's answer
2020-02-20T09:59:07-0500

Answer:


We are given: "f(t) = t^2e^{2t}"


We need to find out the Laplace transform of "f(t)=t^2e^{2t}"


Apply Laplace transformation rule, where "f(t)=e^{2t},k=2"


We need to compute "\frac{d^2}{ds^2}(\frac{1}{s-1})"


Solve by using the chain rule

"\frac{d}{ds}(\frac{1}{s-1})=-\frac{1}{(s-2)^2}\frac{d}{ds}(s-2)=-\frac{1}{(s-2)^2}"


Solve by using the chain rule

"\frac{d^2}{ds^2}(\frac{1}{s-2})=\frac{d^2}{ds^2}(-\frac{1}{(s-2)^2})=\frac{2}{(s-2)^2}\frac{d}{ds}(s-2)=\frac{2}{(s-2)^3}"



Hence "\fbox{L\{$t^2e^{2t}$\}=$\frac{2}{(s-2)^3}$}"